OpenStudy (anonymous):

What is the y-intercept for the graph of this line? 3x - 5y = 6

OpenStudy (diamondboy):

can u isolate y?

OpenStudy (agl202):

The y-intercept is what y is when you substitute 0 in for x. So: 3x - 5y = 15 3(0) - 5y = 15 -5y = 15 y = -3 So the y-intercept is (0, -3).

OpenStudy (k_lynn):

To find the y intercept, all you need to do is plug in 0 for x and solve. So 

OpenStudy (k_lynn):

Right @Agl202 since the x is 0, whatever you get for y is the y coordinate in the ordered pair.

OpenStudy (diamondboy):